OpenStudy (anonymous):

find (dy/dt) when x= -2, dx/dt = 2, x^3 + y^3 = 19

OpenStudy (anonymous):

x^3 + y^3 = 19 3x^2x' + 3y^2y' = 0 -> y' = - x^2x'/ y^2 At x = -2, x' = 2 : y' = - 8/ y^2

OpenStudy (anonymous):

Hey Jinnie, no more related rate, today :)

OpenStudy (anonymous):

Implicit differentiation. Basically we just go ahead and differentiate both sides, but you must remember that you are differentiating with respect to (t), so when you take the derivative of y and x, you must chain rule. \[x^3 + y^3 = 19\] \[d/dt ( x^3 + y^3) = d/dt (19)\] \[3x^2 x' + 3y^2 y' = 0\] now lets go ahead and plug in the numbers they gave us \[3(-2)^2 (-2) + 3y^2 y' = 0\] \[-24 + 3y^2y' = 0\] \[3y^2y' = 24\] \[y' = 24/(3y^2)\] \[y' = 8/(y^2)\] Usually you can just leave the y in the answer, but if you don't want to, just go isolate y in your original equation and plug it in. also remebmer dy/dt is the same as y', just different notation.

OpenStudy (anonymous):

thanks both

OpenStudy (anonymous):

Nada, Jinnie :) @Hermeezey, what a terrific explanation. It's going to cost me the whole day to think like you!

OpenStudy (anonymous):

lol i appreciate both. but is it -8 or +8

OpenStudy (anonymous):

Haha thank you Chlorophyll. I just wanted to be as clear as possible

OpenStudy (anonymous):

It's -8, because when we switch to other side !

OpenStudy (anonymous):

ok thanks. hey cholorophyll can you refresh my memory on what to do to differentiate when you're given (g of f)' (x)

OpenStudy (anonymous):

I always strive for neat and clear, but I'm neither good explainer nor typist :(

OpenStudy (anonymous):

Give me the example, Jinnie!

OpenStudy (anonymous):

 (g of f) (4)